Marston Corporation manufactures disposable thermometers that are sold to hospitals through a network of independent sales agRequired 1. Assuming sales of $39,000,000, construct a budgeted contribution format income statement for the upcoming year fob. The independent sales agents commission rate increases to 20%. (Input all amounts as positive values except losses whichc. The company employs its own sales force. (Input all amounts as positive values except losses which should be indicated by2. Calculate Marston Corporations break-even point in sales dollars for the upcoming year assuming the following: a. The ind3. Refer to your answer to (1)(b) above. If the company employs its own sales force, what volume of sales would be necessary
Answer #1

1
a

Marston Corporation
Budgeted Income Statement
Sales $ 3,90,00,000 100%
Cost of Goods Sold
Variable $ 1,71,00,000
Fixed $      27,20,000
Commission $      70,20,000 $ 2,68,40,000 69%
Contribution $ 1,21,60,000 31%
Fixed Costs
Advertising Expense $        7,20,000
Admin Expense $      32,00,000
Net Income $      82,40,000

b

Marston Corporation
Budgeted Income Statement
Sales $ 3,90,00,000 100%
Cost of Goods Sold
Variable $ 1,71,00,000
Fixed $      27,20,000
Commission $      78,00,000 $ 2,76,20,000 71%
Contribution $ 1,13,80,000 29%
Fixed Costs
Advertising Expense $        7,20,000
Admin Expense $      32,00,000
Net Income $      74,60,000

c

Marston Corporation
Budgeted Income Statement
Sales $ 3,90,00,000 100%
Cost of Goods Sold
Variable $ 1,71,00,000
Fixed $      27,20,000
Commission $      39,00,000 $ 2,37,20,000 61%
Contribution $ 1,52,80,000 39%
Fixed Costs
Sales people Salaries $        6,40,000
Travel and Entertainment $        3,40,000
Sales Manager $        1,90,000
Advertising Expense $      11,60,000
Admin Expense $      32,00,000
Net Income $      97,50,000

2.

a. Breakeven in sales dollar = Fixed Costs / Contribution Margin
= 3,920,000 / 31% = $12,572,368.42

b. Breakeven in sales dollar = Fixed Costs / Contribution Margin
                                             = 3,920,000 / 29% = $13,434,094.90

c. Breakeven in sales dollar = Fixed Costs / Contribution Margin
                                             = 5,530,000 / 39% = $14,114,528.80

3.

Volume of sales necessary = (Fixed Cost + Desired Profit) / Contribution Margin
= (5,530,000 + 7,460,000) / 39%
= $33,307,692

4.

29% of sales - 3,920,000 = 39% of sales - 5,530,000
10% of sales = 5,530,000 - 3,920,000
Sales = $16,100,000